# Ferret: Federated Full-Parameter Tuning at Scale for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have become indispensable in numerous real-world applications. Unfortunately, fine-tuning these models at scale, especially in federated settings where data privacy and communication efficiency are critical, presents significant challenges. Existing methods often resort to parameter-efficient fine-tuning (PEFT) to mitigate communication overhead, but this typically comes at the cost of model accuracy. To address these limitations, we propose federated full-parameter tuning at scale for LLMs (Ferret), the first first-order method with shared randomness to enable scalable full-parameter tuning of LLMs across decentralized data sources while maintaining competitive model accuracy. Ferret accomplishes this through three aspects: (1) it employs widely applied first-order methods for efficient local updates; (2) it projects these updates into a low-dimensional space to considerably reduce communication overhead; and (3) it reconstructs local updates from this low-dimensional space with shared randomness to facilitate effective full-parameter global aggregation, ensuring fast convergence and competitive final performance. Our rigorous theoretical analyses and insights along with extensive experiments, show that Ferret significantly enhances the scalability of existing federated full-parameter tuning approaches by achieving high computational efficiency, reduced communication overhead, and fast convergence, all while maintaining competitive model accuracy. Our implementation is available at https://github.com/allen4747/Ferret.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have become indispensable tools across a wide range of real-world applications, from natural language processing tasks like translation [1] and summarization [2] to more complex tasks such as code generation [3] and decision-making systems [4]. The immense scale and versatility of LLMs make them highly valuable in practice, but they also introduce significant challenges, particularly when they are fine-tuned in federated settings. Federated Learning (FL) offers a decentralized approach to fine-tuning LLMs while retaining data on local clients to ensure privacy. However, while this approach effectively addresses privacy concerns, it also results in prohibitive communication overhead when the model parameters of LLMs scale to billions.
@@
One of the straightforward strategies to mitigate the prohibitive communication costs in the federated tuning of LLMs is parameter-efficient fine-tuning (PEFT). PEFT methods [5, 6] focus on fine-tuning only a subset of model parameters, which is able to significantly reduce the communication overhead between clients and a central server [7, 8, 9, 10]. Despite the effectiveness in reducing bandwidth usage, this type of approach often compromises model accuracy [11], as fine-tuning a subset of model parameters may fail to fully capture the nuances of local data distributions. Thus, recent efforts, e.g., FedKSeed [12], have been devoted to utilizing zeroth-order optimization (ZOO) [13, 14] in federated full-parameter tuning of LLMs, aiming to maintain competitive model accuracy while reducing the communication overhead by transmitting only thousands of scalar gradients per round between clients and a central server. Unfortunately, this approach often suffers from its poor scalability, including increased computational cost per round and a larger number of communication rounds required for convergence, compared to FL methods that use first-order optimization (FOO), e.g., FedAvg [15].
@@
To this end, we propose federated full-parameter tuning at scale for LLMs (Ferret), the first first-order FL method with shared randomness to enable scalable federated full-parameter tuning of LLMs with compelling computational efficiency, reduced communication overhead, and fast convergence speed, all while maintaining competitive model accuracy. Ferret achieves this through three aspects: First, it employs widely applied first-order methods to perform computationally efficient local updates on each client, which typically requires fewer iterations to achieve the same local update process compared to existing ZOO-based FL. Next, Ferret projects these updates into a low-dimensional space, leading to a significantly reduced communication cost compared to existing FOO-based FL. Finally, Ferret reconstructs local updates from the low-dimensional space with shared randomness for effective full-parameter global aggregation, ensuring fast convergence and competitive model accuracy compared to existing ZOO-based FL. We further complement Ferret with rigorous theoretical analyses and principled insights, showing the theoretical advantages of Ferret over other baselines and guiding the best practices for its implementation. Finally, through extensive experiments, we verify that Ferret significantly outperforms existing methods with superior scalability and competitive model accuracy, making it a desirable solution for deploying LLMs in large-scale federated environments.
@@
  • We propose Ferret, the first first-order FL approach with shared randomness (to the best of our knowledge), which significantly enhances the scalability of federated full-parameter tuning of LLMs while maintaining competitive model accuracy.
@@ 2 Problem Setup
In this paper, we consider the federated full-parameter tuning of an LLM using decentralized data $\{{\mathcal{D}}_{i}\}_{i=1}^{\smash{N}}$ on $N$ local clients while preserving data privacy, i.e., without sharing raw data. Specifically, we aim to minimize a global objective ${\mathcal{L}}({\mathbf{w}})$ defined as the average loss across $\{{\mathcal{D}}_{i}\}_{i=1}^{\smash{N}}$ over the model parameters ${\mathbf{w}}\in{\mathbb{R}}^{\smash{d}}$ of an LLM. That is,
$\min_{{\mathbf{w}}}{\mathcal{L}}({\mathbf{w}})\triangleq\frac{1}{N}\sum_{i\in[N]}{\mathcal{L}}^{(i)}({\mathbf{w}})\quad\text{where}\quad{\mathcal{L}}^{(i)}({\mathbf{w}})\triangleq\mathbb{E}_{{\mathbf{x}}^{(i)}\in{\mathcal{D}}_{i}}\left[\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\right]\ .$ (1)
@@
The main challenge in LLM federated full-parameter tuning is to ensure the computational efficiency and the convergence speed of the global model while reducing the communication overheads, particularly given that the parameter size $d$ of LLMs often reaches billions. While existing first-order FL [15, 16, 17] can ensure compelling computational efficiency and convergence speed by applying first-order updates, they typically incur ${\mathcal{O}}(d)$ communication overheads due to the need to transmit the entire set of model parameters between clients and the central server. This type of methods hence is impractical for LLM federated full-parameter tuning due to the enormous size of LLMs. In contrast, although zeroth-order FL [12] can reduce these communication costs by transmitting only several scalar gradients from their finite difference-based gradient estimation with shared randomness, they often incur more computational cost to achieve the same local update progress and a larger number of communication rounds to converge compared with first-order FL. These naturally raise the question: Can we combine the strengths of these methods to achieve scalable federated full-parameter tuning of LLMs with high computational efficiency, reduced communication overhead, and fast convergence?
Input: Pre-trained model parameters ${\mathbf{w}}_{0}$, $N$ clients, number of rounds $R$, number of local updates $T$, number of bases $K$, local learning rate $\eta$
1 for each round $r\in[R]$ do
2       for each client $j\in[N]$ in parallel do
3             if $r>1$ then // Step : Global Aggregation
4                   Receive seeds $\{s^{\smash{(i)}}\}_{i=1}^{N}$ and coordinates $\{\gamma_{k}^{\smash{(i)}}\}_{i=1,k=1}^{N,K}$, generate $\{{\mathbf{v}}_{k}^{\smash{(i)}}\}_{i=1,k=1}^{N,K}$
5                   ${\mathbf{w}}_{r-1}\leftarrow{\mathbf{w}}_{r-2}-\sum_{i\in[N]}\left(\sum_{k=1}^{K}\gamma^{\smash{(i)}}_{k}{\mathbf{v}}^{\smash{(i)}}_{k}\right)/N$
6
7             ${\mathbf{w}}_{r,0}\leftarrow{\mathbf{w}}_{r}$
8            for each iteration $t\in[T]$ do // Step : Local Updates
9                   ${\mathbf{w}}^{\smash{(j)}}_{r,t}\leftarrow{\mathbf{w}}^{\smash{(j)}}_{r,t-1}-\eta\,\nabla\ell({\mathbf{w}}^{\smash{(j)}}_{r,t-1};{\mathbf{x}}_{r,t-1}^{\smash{(j)}})$ on randomly sampled data ${\mathbf{x}}_{r,t-1}^{\smash{(j)}}$
10
             // Step : Projected Updates
11             Randomly choose seeds $s^{\smash{(j)}}$ and generate random bases $\{{\mathbf{v}}_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$
12            $\Delta^{\smash{(j)}}_{r}\leftarrow{\mathbf{w}}^{\smash{(j)}}_{r-1}-{\mathbf{w}}^{\smash{(j)}}_{r}$, compute coordinates $\{\gamma_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ based on (6)
13            Send $s^{\smash{(j)}}$ and $\{\gamma_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ to other clients
Algorithm 1 Federated Full-Parameter Tuning at Scale for LLMs (Ferret)
@@ 3 The Ferret Algorithm
To answer it, we introduce Ferret, federated full-parameter tuning at scale for LLMs, in Algo. 1. We present an overview of our Ferret algorithm in Sec.3.1, followed by a detailed explanation of the key techniques in Ferret in Sec. 3.2.
@@ 3.1 Overview of Ferret
Step 1: Global Aggregation (Line 3-5 in Algo. 1). At the beginning of the first round ($r=1$), each client initializes its local model parameters using the pre-trained model parameters ${\mathbf{w}}_{0}$, i.e., ${\mathbf{w}}_{1}\leftarrow{\mathbf{w}}_{0}$. For subsequent rounds ($r>1$), each client $j\in[N]$ receives the random seeds $s^{\smash{(i)}}$ and the corresponding $K$ projected coordinates $\{\gamma_{k}^{\smash{(i)}}\}_{k=1}^{\smash{K}}$ of every client $i\in[N]$ from the previous round. These random seeds (i.e., shared randomness) are then used to generate $d$-dimensional random bases $\{{\mathbf{v}}_{k}^{\smash{(i)}}\}_{k=1}^{\smash{K}}$ for each client $i$.  [^1] Similar to the practice in [12], we can obtain $K$ random seeds from a single seed $s^{\smash{(i)}}$ and employ these $K$ seeds to generate $K$ random bases independently for each client $i$. So, one seed is sufficient for each client. These random bases, along with the corresponding projected coordinates $\{\gamma_{k}^{\smash{(i)}}\}_{k=1}^{\smash{K}}$, are applied to reconstruct local updates as $\widetilde{\Delta}_{r-1}^{\smash{(i)}}$ in every client $i$. The global model is then updated by aggregating these local contributions as follows:
@@
${\mathbf{w}}_{r-1}\leftarrow{\mathbf{w}}_{r-2}-\frac{1}{N}\sum_{i\in[N]}\widetilde{\Delta}_{r-1}^{\smash{(i)}}\quad\text{where}\quad\widetilde{\Delta}_{r-1}^{\smash{(i)}}\triangleq\sum_{k\in[K]}\gamma_{k}^{(i)}{\mathbf{v}}_{k}^{(i)}\ .$ (2)
@@
Step 2: Local Updates (Line 6-8 in Algo. 1). After Step 1, each client $j$ will perform $T$-iteration first-order optimization on its local loss function by using the randomly sampled data for its local updates. Formally, if stochastic gradient descent with a local learning rate $\eta$ is used, the update rule for client $j\in[N]$ at iteration $t\in[T]$ of round $r\in[R]$ can then be represented as:
${\mathbf{w}}_{r,t}^{(j)}\leftarrow{\mathbf{w}}_{r,t-1}^{(j)}-\eta\nabla\ell\left({\mathbf{w}}_{r,t-1}^{(j)};{\mathbf{x}}_{t-1}^{(j)}\right)\quad\text{where}\quad{\mathbf{w}}_{r,0}\leftarrow{\mathbf{w}}_{r}\ .$ (3)
@@
Different from the zeroth-order update in [12] that requires many local update iterations, the first-order update in (3) enables each client to efficiently and effectively adapt the global model ${\mathbf{w}}_{r}$ to its specific data using a small $T$, thereby enhancing both the computational efficiency of this local update. Here, (3) can be implemented using any gradient method variant, e.g., Adam [18].
@@
Step 3: Projected Updates (Line 9-11 in Algo. 1). After completing the local updates above, each client $j$ randomly chooses a single new seed $s^{(j)}$ to generate $K$ new random bases $\{{\mathbf{v}}_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ and employ these $K$ new random bases to project the local update $\Delta^{\smash{(j)}}_{r}$ into a $K$-dimensional coordinates $\{\gamma_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ based on the techniques in Sec. 3.2. Seed $s^{\smash{(j)}}$ and projected coordinates $\{\gamma_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ are then shared with other clients to facilitate the next round of global aggregation. By sharing only a single random seed and $K$ projected coordinates among $N$ clients where random bases $\{{\mathbf{v}}_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ can be regenerated for global aggregation as shown in Step 1 above, the communication overhead in LLM full-parameter tuning is therefore considerably reduced compared with first-order methods (e.g., FedAvg) especially when $T\ll d$. Of note, the communication of this seed $s^{\smash{(j)}}$ can be mitigated if the same seed is used across all rounds $r\in[R]$, which can further reduce the communication overhead.

### v3

@@ Abstract
Large Language Models (LLMs) have become indispensable in numerous real-world applications. However, fine-tuning these models at scale, especially in federated settings where data privacy and communication efficiency are critical, presents significant challenges. Existing approaches often resort to parameter-efficient fine-tuning (PEFT) to mitigate communication overhead, but this typically comes at the cost of model accuracy. To this end, we propose federated full-parameter tuning at scale for LLMs (Ferret), the first first-order method with shared randomness to enable scalable full-parameter tuning of LLMs across decentralized data sources while maintaining competitive model accuracy. Ferret accomplishes this through three aspects: (I) it employs widely used first-order methods for efficient local updates; (II) it projects these updates into a low-dimensional space to considerably reduce communication overhead; and (III) it reconstructs local updates from this low-dimensional space with shared randomness to facilitate effective full-parameter global aggregation, ensuring fast convergence and competitive final performance. Our rigorous theoretical analyses and insights along with extensive experiments, show that Ferret significantly enhances the scalability of existing federated full-parameter tuning approaches by achieving high computational efficiency, reduced communication overhead, and fast convergence, all while maintaining competitive model accuracy. Our implementation is available at https://github.com/allen4747/Ferret.
Machine Learning, ICML
Figure 1: Performance comparison of various federated full-parameter tuning algorithms on Natural Instructions dataset with LLaMA-3B. Our Ferret shows significantly improved scalability, with a 6.0$\times$ reduction in computational cost and 3.3$\times$ fewer convergence rounds than FedKSeed, alongside a $\text{10}^{\text{6}}\times$ reduction in communication overhead than FedAvg, while achieving comparable test score.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have become indispensable tools across a wide range of real-world applications, from natural language processing tasks like translation (Xu et al., 2024) and summarization (Van Veen et al., 2024) to more complex tasks such as code generation (Liu et al., 2024) and decision-making systems (Shao et al., 2023). The immense scale and versatility of LLMs make them highly valuable in practice, but they also introduce significant challenges, particularly when they are fine-tuned in federated settings. Federated Learning (FL) offers a decentralized approach to fine-tuning LLMs while retaining data on local clients to ensure privacy. However, while this approach effectively addresses privacy concerns, it also results in prohibitive communication overhead when the model parameters of LLMs scale to billions.
@@
One of the straightforward strategies to mitigate the prohibitive communication costs in the federated tuning of LLMs is parameter-efficient fine-tuning (PEFT). PEFT methods (Hu et al., 2022; Wei et al., 2024) focus on fine-tuning only a subset of model parameters, which is able to significantly reduce the communication overhead between clients and a central server (Che et al., 2023; Zhang et al., 2023; Kuang et al., 2024; Zhang et al., 2024b). Despite the effectiveness in reducing bandwidth usage, this type of approach often compromises model accuracy (Pu et al., 2023), as fine-tuning a subset of model parameters may fail to fully capture the nuances of local data distributions. Thus, recent efforts, e.g., FedKSeed (Qin et al., 2024), have been devoted to utilizing zeroth-order optimization (ZOO) (Nesterov & Spokoiny, 2017; Berahas et al., 2022) in federated full-parameter tuning of LLMs, aiming to maintain competitive model accuracy while reducing the communication overhead by transmitting only thousands of scalar gradients per round between clients and a central server. Unfortunately, this approach often suffers from its poor scalability, including increased computational cost per round and a larger number of communication rounds required for convergence, compared to FL methods that use first-order optimization (FOO), e.g., FedAvg (McMahan et al., 2017).
@@
Therefore, we develop federated full-parameter tuning at scale for LLMs (Ferret), the first first-order FL approach with shared randomness to enable scalable federated full-parameter tuning of LLMs with compelling computational efficiency, reduced communication overhead, and fast convergence speed, while maintaining competitive model accuracy, as shown in Fig. 1. Ferret achieves this through three aspects: First, it uses widely applied first-order methods to perform efficient local updates on each client, which usually requires fewer iterations to achieve the same local update process compared to existing ZOO-based FL. Next, Ferret projects these updates into a low-dimensional space, resulting in a significantly reduced communication cost compared to existing FOO-based FL. Finally, Ferret reconstructs local updates from the low-dimensional space with shared randomness for effective full-parameter global aggregation, ensuring fast convergence and competitive model accuracy compared to existing ZOO-based FL. We further complement Ferret with rigorous theoretical analyses and insights, showing the theoretical advantages of Ferret over other baselines and guiding the best practices for its implementation. Finally, through extensive experiments, we verify that Ferret significantly outperforms existing methods with superior scalability and competitive model accuracy, making it a desirable solution for deploying LLMs in large-scale federated environments.
@@
  • $\bullet$
    We novelly propose Ferret, to the best of our knowledge, the first first-order FL approach with shared randomness, which significantly enhances the scalability of federated full-parameter tuning of LLMs while maintaining competitive model accuracy.
  • $\bullet$
@@ 2 Problem Setup
In this paper, we consider the federated full-parameter tuning of an LLM using decentralized data $\{{\mathcal{D}}_{i}\}_{i=1}^{\smash{N}}$ on $N$ local clients while preserving data privacy, i.e., without sharing raw data. Specifically, given a loss function $\ell(\cdot;\cdot)$, we aim to minimize a global objective ${\mathcal{L}}({\mathbf{w}})$ defined as the average loss across $\{{\mathcal{D}}_{i}\}_{i=1}^{\smash{N}}$ over the model parameters ${\mathbf{w}}\in{\mathbb{R}}^{\smash{d}}$ of an LLM. That is,
$\min_{{\mathbf{w}}}{\mathcal{L}}({\mathbf{w}})$ $\triangleq\frac{1}{N}\sum_{i\in[N]}{\mathcal{L}}^{(i)}({\mathbf{w}})$ (1)
$\text{where}\quad{\mathcal{L}}^{(i)}({\mathbf{w}})$ $\triangleq\mathbb{E}_{{\mathbf{x}}^{(i)}\in{\mathcal{D}}_{i}}\left[\ell({\mathbf{w}};{\mathbf{x}}^{(i)})\right]\ .$
@@
The main challenge in LLM federated full-parameter tuning is to ensure the computational efficiency and the convergence speed of the global model while reducing the communication overheads, particularly given that the parameter size $d$ of LLMs often reaches billions. While existing first-order FL (McMahan et al., 2017; Li et al., 2020; Karimireddy et al., 2020) can ensure compelling computational efficiency and convergence speed by applying first-order updates, they typically incur ${\mathcal{O}}(d)$ communication overheads due to the need to transmit the entire set of model parameters between clients and the central server. This type of methods hence is impractical for LLM federated full-parameter tuning due to the enormous size of LLMs. In contrast, although zeroth-order FL (Qin et al., 2024) can reduce these communication costs by transmitting only several scalar gradients from their finite difference-based gradient estimation with shared randomness, they often incur more computational cost to achieve the same local update progress and a larger number of communication rounds to converge compared with first-order FL. These naturally raise the question:
Can we combine the strengths of these different types of methods to achieve scalable federated full-parameter tuning of LLMs with high computational efficiency, reduced communication overhead, and fast convergence?
@@ 3 The Ferret Algorithm
To answer this question, we introduce Ferret, federated full-parameter tuning at scale for LLMs, in Algo. 1. We present an overview of Ferret algorithm in Sec.3.1, followed by a detailed explanation of its key techniques in Sec. 3.2.
@@ 3.1 Overview of Ferret
Step 1: Global Aggregation (Line 3-6 in Algo. 1). At the beginning of the first round ($r=1$), each client initializes its local model parameters using the pre-trained model parameters ${\mathbf{w}}_{0}$, i.e., ${\mathbf{w}}_{1}{\leftarrow}{\mathbf{w}}_{0}$. For subsequent rounds ($r>1$), each client $j\in[N]$ receives the random seeds $s^{\smash{(i)}}$ and the corresponding $K$ projected coordinates $\{\gamma_{k}^{\smash{(i)}}\}_{k=1}^{\smash{K}}$ of every client $i\in[N]$ from the previous round. These random seeds (i.e., shared randomness) are then used to generate $d$-dimensional random bases $\{{\mathbf{v}}_{k}^{\smash{(i)}}\}_{k=1}^{\smash{K}}$ for each client $i$.  [^1] As in (Qin et al., 2024), we can get $K$ random seeds from a single seed $s^{\smash{(i)}}$ and use them to generate $K$ random bases independently. So, one seed is sufficient for each client. These random bases, along with the corresponding projected coordinates $\{\gamma_{k}^{\smash{(i)}}\}_{k=1}^{\smash{K}}$, are applied to reconstruct local updates as $\widetilde{\Delta}_{r-1}^{\smash{(i)}}$ in every client $i$. The global model is then updated by aggregating these local contributions as follows:
@@
${\mathbf{w}}_{r-1}{\leftarrow}{\mathbf{w}}_{r-2}{-}\frac{1}{N}\sum_{i\in[N]}\widetilde{\Delta}_{r-1}^{\smash{(i)}},\text{with}\ \widetilde{\Delta}_{r-1}^{\smash{(i)}}{\triangleq}\sum_{k\in[K]}\gamma_{k}^{(i)}{\mathbf{v}}_{k}^{(i)}\ .$ (2)
@@
Step 2: Local Updates (Line 7-9 in Algo. 1). After Step 1, each client $j$ will perform $T$-iteration first-order optimization on its local loss function by using the randomly sampled data for its local updates. Formally, if stochastic gradient descent with a local learning rate $\eta$ is used, the update rule for client $j\in[N]$ at iteration $t\in[T]$ of round $r\in[R]$ can then be represented as below with ${\mathbf{w}}_{r,0}\leftarrow{\mathbf{w}}_{r}$:
${\mathbf{w}}_{r,t}^{(j)}{\leftarrow}{\mathbf{w}}_{r,t-1}^{(j)}{-}\eta\nabla\ell\left({\mathbf{w}}_{r,t-1}^{(j)};{\mathbf{x}}_{t-1}^{(j)}\right)\ .$ (3)
@@
Different from the zeroth-order update in (Qin et al., 2024) that requires many local update iterations, the first-order update in (3) enables each client to efficiently and effectively adapt the global model ${\mathbf{w}}_{r}$ to its specific data using a small $T$, thereby enhancing both the computational efficiency of this local update. Here, (3) can be implemented using any gradient method variant, e.g., Adam (Kingma & Ba, 2014).
@@
Step 3: Projected Updates (Line 10-12 in Algo. 1). After completing the local updates above, each client $j$ randomly chooses a single new seed $s^{(j)}$ to generate $K$ new random bases $\{{\mathbf{v}}_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ and employ these $K$ new random bases to project the local update $\Delta^{\smash{(j)}}_{r}$ into a $K$-dimensional coordinates $\{\gamma_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ based on the techniques in Sec. 3.2. Seed $s^{\smash{(j)}}$ and projected coordinates $\{\gamma_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ are then shared with other clients to facilitate the next round of global aggregation. By sharing only one single random seed and $K$ projected coordinates among $N$ clients where random bases $\{{\mathbf{v}}_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ can be regenerated for global aggregation as shown in Step 1 above, the communication overhead in LLM full-parameter tuning is therefore considerably reduced compared with first-order methods (e.g., FedAvg) especially when $T\ll d$. The communication of seed $s^{\smash{(j)}}$ can be mitigated if the same seed is used across all rounds $r\in[R]$, which can further reduce communication overhead.
Algorithm 1 Ferret
Input: ${\mathbf{w}}_{0}$, $N$, $R$, $T$, $K$, $\eta$
for each round $r\in[R]$ do
2       for each client $j\in[N]$ in parallel do
3             if $r>1$ then // Step : Global Aggregation
4                   Receive $\{s^{\smash{(i)}}\}_{i=1}^{N}$ and $\{\gamma_{k}^{\smash{(i)}}\}_{i=1,k=1}^{N,K}$  Generate bases $\{{\mathbf{v}}_{k}^{\smash{(i)}}\}_{i=1,k=1}^{N,K}$ using $\{s^{\smash{(i)}}\}_{i=1}^{N}$${\mathbf{w}}_{r-1}{\leftarrow}{\mathbf{w}}_{r-2}{-}\sum_{i\in[N]}\left(\sum_{k=1}^{K}\gamma^{\smash{(i)}}_{k}{\mathbf{v}}^{\smash{(i)}}_{k}\right)/N$
5             ${\mathbf{w}}_{r,0}\leftarrow{\mathbf{w}}_{r}$ for $t\in[T]$ do // Step : Local Updates
6                   ${\mathbf{w}}^{\smash{(j)}}_{r,t}\leftarrow{\mathbf{w}}^{\smash{(j)}}_{r,t-1}-\eta\,\nabla\ell({\mathbf{w}}^{\smash{(j)}}_{r,t-1};{\mathbf{x}}_{r,t-1}^{\smash{(j)}})$
7             // Step : Projected Updates Randomly set $s^{\smash{(j)}}$ and generate bases $\{{\mathbf{v}}_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ $\Delta^{\smash{(j)}}_{r}{\leftarrow}{\mathbf{w}}^{\smash{(j)}}_{r-1}{-}{\mathbf{w}}^{\smash{(j)}}_{r}$, compute $\{\gamma_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ with (6) Send $s^{\smash{(j)}}$ and $\{\gamma_{k}^{\smash{(j)}}\}_{k=1}^{\smash{K}}$ to the central server